_______
 Ö8x - 17 - 1 = 0

Answer by AnlytcPhil(1810)


  _______
 Ö8x - 17 - 1 = 0

Isolate the radical term

      _______
     Ö8x - 17 = 1

Square both sides:
    _______
  (Ö8x - 17)² = (1)²
 
      8x - 17 = 1

          8x = 18

           x = 18/8

           x = 9/4

We must check a radical equation because sometimes
an answer we get does not check even though we did
nothing wrong.

       _______
      Ö8x - 17 - 1 = 0 
   ___________
  Ö8(9/4) - 17 - 1 = 0
       _______
      Ö18 - 17 - 1 = 0
            _
           Ö1 - 1 = 0
            
            1 - 1 = 0

                0 = 0

So x = 9/4 checks and thus is the only solution.

Edwin